# Towards Fair and Efficient Public Transportation: A Bus Stop Model

### v1

@@ Abstract
We consider a stylized formal model of public transportation, where a set of agents need to travel along a given road, and there is a bus that runs the length of this road. Each agent has a left terminal and a right terminal between which they wish to travel; they can walk all the way, or walk to/from the nearest stop and use the bus for the rest of their journey. The bus can make a fixed number of stops, and the planner needs to select locations for these stops. We study notions of efficiency and fairness for this setting. First, we give a polynomial-time algorithm for computing a solution that minimizes the total travel time; our approach can capture further extensions of the base model, such as more general cost functions or existing infrastructure. Second, we develop a polynomial-time algorithm that outputs solutions with provable fairness guarantees (such as a variant of the justified representation axiom or $2$-approximate core) as long as the agents’ costs only depend on the distance they need to walk. Our simulations indicate that our algorithm almost always outputs fair solutions, even for parameter regimes that do not admit theoretical guarantees.
@@ 1 Introduction
The need to model and solve problems related to public transport has been under scrutiny from an operations research perspective; see, e.g., the extensive survey by Desaulniers and Hickman (2007). In the optimization literature, the implementation of public transport infrastructure is commonly seen as a two-stage process consisting of a planning phase and an operational phase. The planning phase is concerned with the design of the transportation network as well as with determining optimal operation frequencies (Lampkin and Saalmans, 1967; Silman et al., 1974). In the operational phase, the cost of operating public transport should be minimized, e.g., by optimally assigning vehicles to routes or drivers to buses (Daduna and Pinto Paixão, 1995; Wren and Rousseau, 1995). In both phases, the primary metric used to evaluate the solution quality is the social welfare, i.e., the total/average travel time.
@@
We propose to tackle this challenge using the conceptual apparatus of group fairness, building on the ideas of justified representation in multi-winner voting (Aziz et al., 2017) and core stability in cooperative game theory (Gillies, 1959). The intuition that we aim to capture is that sufficiently large groups of agents with similar preferences deserve to be represented in the selected solution, or, more ambitiously, that each group should be allocated resources in proportion to its size.
While we believe that this perspective should be taken into account at all stages of transportation planning, we showcase our approach by applying it to a specific and relatively simple task: choosing the locations of the stops for a fixed bus/train route. Specifically, we consider the setting where the trajectory of the vehicle has been exogenously determined, either by topography (e.g., a mountain road or a river) or by existing infrastructure (e.g., train tracks), the number of stops has been fixed in advance due to bounds on the overall travel time, but the designer still has the freedom to decide where to place the stops. Then, to use the public transport option, the user would have to travel to a nearby stop by using private transport (such as walking, cycling, or using an e-scooter), ride the bus towards their destination and then use private transport again for the last-mile travel. Alternatively, they can opt to use private transport for the entire trip; however, we assume that private transport has higher per-mile cost (measured as physical effort, travel time, or monetary cost) than public transport. Crucially, the agents’ decision whether to use the public transport at all is influenced by the location of the stops, so the planner’s choices made at this stage may have a dramatic effect on the demand for public transport: positioning the stops without taking into account the agents’ travel needs may render the system unusable and push the residents towards private transportation solutions.
@@ 1.1 Our Contribution
[…]
In the second part of the paper, we focus on finding solutions that provide justified representation (JR) or are (approximately) in the core. Unfortunately, efficiency and justified representation turn out to be incompatible. However, we present a polynomial-time algorithm that operates by selecting bus stops at distances proportional to the density of terminal points, and show that this algorithm finds JR solutions whenever the cost of taking the bus is zero. In fact, this algorithm offers a $2$-approximation to the stronger fairness concept of the core, and exhibits excellent empirical performance (on synthetic data). In contrast, there are instances for which no solution can provide strong JR.
@@ 1.2 Related Work
In contrast, our approach, i.e., modeling fairness in terms of proportionality, is rooted in the (computational) social choice literature (see e.g., Conitzer et al., 2017; Jiang et al., 2020; Peters and Skowron, 2020). Our model can be viewed as a special case of multi-winner voting, and our notion of justified representation is an adaptation of a similar concept in multi-winner approval voting (Aziz et al., 2017). It is also similar to the notion of proportional fairness in fair clustering (Chen et al., 2019; Micha and Shah, 2020; Aziz et al., 2024a). In this stream of literature, Li et al. (2021) study approximate core stability, where the approximation is with respect to the size of the deviating coalition (which is similar in spirit to our approach) or with respect to the gain by the deviating agents. Kalayci et al. (2024) consider a similar approximation in the context of multi-winner voting, and Chaudhury et al. (2022) explore similar ideas in the context of federated learning.
We note that placing stops on the line is similar in spirit to facility location (Chan et al., 2021); however, our focus in this work is on fairness, whereas much of the facility location literature takes a mechanism design perspective (see, however, Zhou et al. (2022); Elkind et al. (2022)). Most related to our paper are models which investigate the same cost function (Fukui et al., 2020; Chan and Wang, 2023). In particular, the model by Chan and Wang (2023) is a special cost of our model where $\alpha=0$ (i.e., taking the bus has no cost), all agents have the same destination, and only two bus stops are built. However, our work differs in two key aspects: we allow for more than two stops to be build and study fairness aspects (rather than strategic manipulation). The facility location literature also considers agents that are interested in more than one location (Serafino and Ventre, 2014; Anastasiadis and Deligkas, 2018), but these works use different cost functions.
@@ 2 Model
Given a positive integer $k\in\mathbb{N}$, we write $[k]:=\{1,\dots,k\}$. For two numbers $x,y\in\mathbb{Q}$, we denote by $d(x,y):=|x-y|$ the Euclidean distance from $x$ to $y$. We extend this notation to sets: given a number $x\in\mathbb{Q}$ and a set of numbers $P\subseteq\mathbb{Q}$, we write $d(x,P):=\min_{y\in P}d(x,y)$.
@@
$c_{i}^{\mathcal{I}}(S):=\min\begin{cases}d(\ell_{i},r_{i})\\ \min_{x,y\in S}\left[d(\ell_{i},x)+\alpha\cdot d(x,y)+d(r_{i},y)\right]\text{.}\end{cases}$ (1)
@@
Apart from efficiency, we are also interested in fairness. Our first concept of fairness builds on ideas from the multi-winner voting and fair clustering literature (Aziz et al., 2017; Chen et al., 2019; Micha and Shah, 2020; Li et al., 2021; Aziz et al., 2024b). Suppose we are given an instance with $n$ agents and budget $b$. Then, intuitively, each agent is entitled to $\frac{b}{n}$ units of money, so a group of $\lceil\frac{n}{b}\rceil$ agents should be able to dictate the position of one stop. Therefore, one may want to rule out solutions $S$ such that all agents in a group of size at least $\frac{n}{b}$ can lower their costs by abandoning $S$ and building a single stop. However, this condition is too weak, as no agent benefits from a single stop. Hence, we strengthen it by considering groups of agents that are entitled to two stops.
@@
A solution $S\subseteq V$ is said to provide justified representation (JR) if, for every set of agents $M\subseteq N$ with $|M|\geq\frac{2n}{b}$ and every pair of stops $T\subseteq V$ there exists an agent $i\in M$ such that $c_{i}(T)\geq c_{i}(S)$. Moreover, a solution $S\subseteq V$ is said to provide strong justified representation if, for every set of agents $M\subseteq N$ with $|M|\geq\frac{2n}{b}$ and every pair of stops $T\subseteq V$ there exists an agent $i\in M$ such that $c_{i}(T)>c_{i}(S)$ or for all agents $i\in M$ it holds that $c_{i}(T)\geq c_{i}(S)$.
The key distinction between JR and strong JR is that, to define the former, we only consider deviations to pairs of stops that are strictly better for each agent in $M$, whereas to define the latter, we also consider deviations that make no agent in $M$ worse off while making at least one member of $M$ strictly better off. Thus, an outcome that provides strong JR also provides JR, but the converse is not necessarily true.
@@
We now show every other solution costs more than $S^{*}$. Fix a solution $S\subseteq V$ with $|S|=6$. If $1\notin S$, then the walking cost of each of the agents in $\{a_{3},a_{4},a_{5},a_{6}\}$ is at least $1$, so $c(S)\geq 60\alpha+4(1-\alpha)$. Therefore, we may assume that $1\in S$. Moreover, if $15\notin S$, then the walking costs of $a_{1}$ and $a_{2}$ are at least $2$, so the solution is worse than $S^{*}$. Hence, we may also assume that $15\in S$.
@@
In fact, the incompatibility observed in Proposition 2.5 can be strengthened further: it holds for approximate JR and, in case of the 00-BSP, even for approximate minimum cost. We provide the details for these results in Appendix A. However, if we replace cost minimality with Pareto optimality, the incompatibility no longer holds: by applying Pareto improvements, we can transform a solution providing JR into a Pareto-optimal solution providing JR.
@@ 3 Efficiency
Our algorithm is based on a dynamic program, which iteratively considers adding new stops to the solution. Capturing our problem by a dynamic program is challenging, because each agent’s cost depends on the placement of two stops. The crucial observation that enables us to circumvent this difficulty is that, to perform cost updates in the dynamic program, it suffices to know the rightmost stop in the current solution. This idea is formalized by Lemma 3.1. All missing proofs from this section can be found in Appendix B.
@@
For $\alpha\in[0,1]$, we can compute a minimum-cost solution for an instance of $\alpha$-BSP in time $O(nm+m^{3})$.
@@
Let $\alpha\in[0,1]$, and consider an instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of $\alpha$-BSP where $V=\{v_{1},\dots,v_{m}\}$ with $v_{1}\leq v_{2}\leq\dots\leq v_{m}$, i.e., potential stops are sorted from left to right and $v_{j}$ represents the $j$-th stop. We will assume $b\leq m$, as otherwise there is an optimal-cost solution that builds a stop at each location. For a solution $S$ with $\max S=v_{j}$ and $k>j$, let $\Delta(j,k):=c(S)-c(S\cup\{v_{k}\})$ denote the total reduction in the agents’ costs from adding stop $v_{k}$ to $S$. By Lemma 3.1, we know that $\Delta(j,k)$ only depends on $v_{k}$ and $v_{j}$, and can be computed in $O(1)$ time. Let us set up a dynamic program $\mathbf{dp}[h,c]$, where
  • $h\in\{0,1,\dots,m\}$ represents the rightmost stop that has been added to the solution, where 00 means that no stop has been added yet, and
@@
That is, we consider the position of the stop $h^{\prime}$ that precedes $h$ in the solution, and evaluate the cost reduction from adding $h$ to a solution that end with $h^{\prime}$. Clearly, the updates can be computed in time $O(m)$, using the update formulas provided by Lemma 3.1. As we assume that $b\leq m$, our table has at most $m^{2}$ entries, and can be filled in time $O(m^{3})$. Hence, the total running time is $O(nm+m^{3})$.
It is not hard to see that our dynamic program is correct; for completeness, we provide the proof of the following claim in Appendix B.
Claim 3.3.
Hence, the minimum cost of a feasible solution for $\mathcal{I}$ is equal to $\min_{h\in\{0,\dots,m\}}\mathbf{dp}[h,b]\text{.}$ We can efficiently extract an explicit feasible solution of minimum cost by standard techniques. ∎
We note that we can also assume that $b\leq 2n$ (otherwise, we can build a stop at each terminal), and hence the running time of our algorithm can also be bounded as $O(n^{2}+nm^{2})$. However, we expect this bound to be less useful than the one stated in Theorem 3.2, since in practical applications it is likely that $m<<n$.
@@
First, the theorem extends to more general cost functions. Consider a setting where the time to travel between two stops depends on factors other than the distance. For instance, the bus route might encompass intervals with different speed limits, or there may be hilly or curvy roads, where the bus needs to slow down. Hence, the costs for travelling need not be homogeneous. However, typically the travel cost of a route only depends on the costs of its segments. To capture this, we introduce the notion of separable travel costs, and formally introduce the separable-cost BSP problem, which generalizes the $\alpha$-BSP problem defined earlier in the paper.
A function $d\colon\mathbb{Q}^{2}\to\mathbb{Q}$ is called separable if for all $x<y<z$ it holds that $d(x,z)=d(x,y)+d(y,z)$. An instance $\mathcal{I}$ of separable-cost BSP is given by a tuple $\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ and separable cost functions $d^{B}:V\times V\to{\mathbb{Q}}$ and $d_{i}^{W}:V\times V\to{\mathbb{Q}}$ for all $i\in N$ (where the superscripts refer to walking and taking the bus). The agents’ costs are then defined as
@@
For separable-cost BSP, we can compute the minimum cost in time $O(nm+m^{3})$.
@@
For $\alpha$-BSP with existing bus stops, we can compute the minimum cost in time $O(nm+m^{3})$.
As a third extension, we consider the case where bus stops do not have identical costs: indeed, construction costs may vary depending on, e.g., ease of access. An instance of $\alpha$-BSP with bus stop costs consists of and instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of the base model together with a budget function $\gamma\colon V\to\mathbb{N}$. A solution $S\subseteq V$ is feasible if $\sum_{i\in S}\gamma(i)\leq b$. Clearly, we can still apply the dynamic program in Theorem 3.2. However, we can no longer assume that $b\leq m$. Hence, we obtain a running time of $O(nm^{2}+m^{3}b)$, which is only pseudo-polynomial due to the dependency on $b$. If the bus stop costs are represented by unary numbers, the running time remains polynomial. However, for bus stop costs represented by binary numbers, we obtain a computational hardness result, via a reduction from Knapsack.
@@
[…]
We conclude this section with a structural result regarding minimum-cost solutions. Interestingly, as long as both terminals of all agents belong to the set of potential bus stops, there is a minimum-cost solution that places all stops at the agents’ terminals. We remark that this is the only place in this section where we need the assumption $\mathcal{A}(\mathcal{I})\subseteq V$.

### v2

@@ Abstract
We consider a stylized formal model of public transportation, where a set of agents need to travel along a given road, and there is a bus that runs the length of this road. Each agent has a left terminal and a right terminal between which they wish to travel; they can walk all the way, or walk to/from the nearest stop and use the bus for the rest of their journey. The bus can make a fixed number of stops, and the planner needs to select locations for these stops. We study notions of efficiency and fairness for this setting. First, we give a polynomial-time algorithm for computing a solution that minimizes the total travel time; our approach can capture further extensions of the base model, such as more general cost functions or existing infrastructure. Second, we develop a polynomial-time algorithm that outputs solutions with provable fairness guarantees (such as a variant of the justified representation axiom or $2$-approximate core). Our simulations indicate that our algorithm almost always outputs fair solutions, even for parameter regimes that do not admit theoretical guarantees.
@@ 1 Introduction
The need to model and solve problems related to public transport has been under scrutiny from an operations research perspective; see, e.g., Desaulniers and Hickman (2007) for an extensive literature survey. In the optimization literature, the implementation of public transport infrastructure is commonly seen as a two-stage process consisting of a planning phase and an operational phase. The planning phase is concerned with the design of the transportation network as well as with determining optimal operation frequencies (Lampkin and Saalmans, 1967; Silman et al., 1974). In the operational phase, the cost of operating public transport should be minimized, e.g., by optimally assigning vehicles to routes or drivers to buses (Daduna and Pinto Paixão, 1995; Wren and Rousseau, 1995). In both phases, the primary metric used to evaluate the solution quality is the social welfare, i.e., the total/average travel time.
@@
We propose to tackle this challenge using the conceptual apparatus of group fairness, building on the ideas of justified representation in multi-winner voting Aziz et al. (2017) and core stability in cooperative game theory Gillies (1959). The intuition that we aim to capture is that sufficiently large groups of agents with similar preferences deserve to be represented in the selected solution, or, more ambitiously, that each group should be allocated resources in proportion to its size.
While we believe that this perspective should be taken into account at all stages of transportation planning, we showcase our approach by applying it to a specific and relatively simple task: choosing the locations of the stops for a fixed bus/train route. Specifically, we consider the setting where the trajectory of the vehicle has been exogenously determined, either by topography (e.g., a mountain road or a river) or by existing infrastructure (e.g., train tracks), the number of stops has been fixed in advance due to bounds on the overall travel time, but the designer still has the freedom to decide where to place the stops. Then, to use the public transport option, the user would have to travel to a nearby stop by using private transport (such as walking, cycling, or using an e-scooter), ride the vehicle towards their destination, and then use private transport again for the last-mile travel. Alternatively, they can opt to use private transport for the entire trip; however, we assume that private transport has higher per-mile cost (measured as physical effort, travel time, or monetary cost) than public transport. Crucially, the agents’ decision whether to use the public transport at all is influenced by the location of the stops, so the planner’s choices made at this stage may have a dramatic effect on the demand for public transport: positioning the stops without taking into account the agents’ travel needs may render the system unusable and push the residents towards private transportation solutions.
@@ 1.1 Our Contribution
Our first contribution is a dynamic program that can efficiently compute cost-minimal solutions. This approach is very flexible in that it still works when we add further features to the model, such as travel costs dependent on non-homogeneous road conditions or existing infrastructure. Moreover, while computing the minimum total cost becomes $\mathsf{NP}$-complete when bus stops have variable costs, our dynamic program still runs in pseudo-polynomial time with respect to the budget.
In the second part of the paper, we focus on finding solutions that provide justified representation (JR) or are (approximately) in the core. Unfortunately, efficiency and justified representation turn out to be incompatible. However, we present a polynomial-time algorithm that operates by selecting bus stops at distances proportional to the density of terminal points, and show that this algorithm finds JR solutions whenever the cost of taking the bus is zero. Moreover, this algorithm offers a $2$-approximation to the stronger fairness concept of the core, and exhibits excellent empirical performance (on synthetic data). In contrast, there are instances for which no solution can provide a stronger form of JR.
@@ 1.2 Related Work
In contrast, our approach, i.e., modeling fairness in terms of proportionality, is rooted in the (computational) social choice literature (see e.g., Conitzer et al., 2017; Jiang et al., 2020; Peters and Skowron, 2020). Our model can be viewed as a special case of multi-winner voting, and our notion of justified representation is an adaptation of a similar concept in multi-winner approval voting Aziz et al. (2017). It is also similar to the notion of proportional fairness in fair clustering (Chen et al., 2019; Micha and Shah, 2020). In this stream of literature, Li et al. (2021) study approximate core stability, where the approximation is with respect to the size of the deviating coalition (which is similar in spirit to our approach) or with respect to the gain by the deviating agents. Kalayci et al. (2024) consider a similar approximation in the context of multi-winner voting, and Chaudhury et al. (2022) explore similar ideas in federated learning settings.
We note that placing stops on the line is similar in spirit to facility location Chan et al. (2021); however, our focus in this work is on fairness, whereas much of the facility location literature takes a mechanism design perspective (see, however, Zhou et al. (2022); Elkind et al. (2022)). Most related to our paper are models which investigate the same cost function Fukui et al. (2020); Chan and Wang (2023). In particular, the model by Chan and Wang (2023) is a special case of our model where $\alpha=0$ (i.e., taking the bus has no cost), all agents have the same destination, and only two bus stops are built. However, our work differs in two key aspects: we allow for more than two stops to be built and study fairness aspects (rather than strategic manipulation). The facility location literature also considers agents that are interested in more than one location Serafino and Ventre (2014); Anastasiadis and Deligkas (2018), but these works use different cost functions.
@@ 2 Model
Given a positive integer $k\in\mathbb{N}$, we write $k]:=\{1,\dots,k\}$. For two numbers $x,y\in\mathbb{Q}$, we denote by d(x,y):=|xy|d(x,y):=|x-y| the Euclidean distance from $x$ to $y$. We extend this notation to sets: given a number $x\in\mathbb{Q}$ and a set of numbers $P\subseteq\mathbb{Q}$, we write $d(x,P):=\min_{y\in P}d(x,y)$.
@@
$c_{i}^{\mathcal{I}}(S):=\min\begin{cases}d(\ell_{i},r_{i})&\text{``agent walks''}\\ \min_{x,y\in S}\left[d(\ell_{i},x)+\alpha\cdot d(x,y)+d(r_{i},y)\right]&\text{``agent takes bus''}.\end{cases}$ (1)
@@
[…]
Figure 1: Illustration of Example 2.4. The same instance proves the incompatibility of efficiency and JR in Proposition 2.5.
@@
Apart from efficiency, we are also interested in fairness. Our first concept of fairness builds on ideas from the multi-winner voting and fair clustering literature (Aziz et al., 2017; Chen et al., 2019; Micha and Shah, 2020; Li et al., 2021; Aziz et al., 2024). Suppose we are given an instance with $n$ agents and budget $b$. Then, intuitively, each agent is entitled to $\frac{b}{n}$ units of money, so a group of $\lceil\frac{n}{b}\rceil$ agents should be able to dictate the position of one stop. Therefore, one may want to rule out solutions $S$ such that all agents in a group of size at least $\frac{n}{b}$ can lower their costs by abandoning $S$ and building a single stop. However, this condition is too weak, as no agent benefits from a single stop. Hence, we strengthen it by considering groups of agents that are entitled to two stops.
@@
A solution $S\subseteq V$ is said to provide justified representation (JR) if for every set of agents $M\subseteq N$ with $|M|\geq\frac{2n}{b}$ and every pair of stops $T\subseteq V$ there exists an agent $i\in M$ such that $c_{i}(T)\geq c_{i}(S)$. Moreover, a solution $S\subseteq V$ is said to provide strong justified representation if for every set of agents $M\subseteq N$ with $|M|\geq\frac{2n}{b}$ and every pair of stops $T\subseteq V$ there exists an agent $i\in M$ such that $c_{i}(T)>c_{i}(S)$ or for all agents $i\in M$ it holds that $c_{i}(T)\geq c_{i}(S)$.
The key distinction between JR and strong JR is that, to define the former, we only consider deviations to pairs of stops that are strictly preferred by each agent in $M$, whereas to define the latter, we also consider deviations that make no agent in $M$ worse off while making at least one member of $M$ strictly better off. Thus, an outcome that provides strong JR also provides JR, but the converse is not necessarily true.
@@
We now show that every other solution costs more than $S^{*}$. Fix a solution $S\subseteq V$ with $|S|=6$. If $1\notin S$, then the walking cost of each of the agents in $\{a_{3},a_{4},a_{5},a_{6}\}$ is at least $1$, so $c(S)\geq 60\alpha+4(1-\alpha)$. Therefore, we may assume that $1\in S$. Moreover, if $15\notin S$, then the walking costs of $a_{1}$ and $a_{2}$ are at least $2$, so the solution is worse than $S^{*}$. Hence, we may also assume that $15\in S$.
@@
In fact, the incompatibility observed in Proposition 2.5 can be strengthened further: it holds for approximate JR and, in case of the 0-BSP, even for approximate minimum cost. We provide the details for these results in Appendix A. However, if we replace cost minimality with Pareto optimality, the incompatibility no longer holds: by applying Pareto improvements, we can transform a solution providing JR into a Pareto-optimal solution providing JR.
Definition 2.6.
Proposition 2.7.
Proof.
Proposition 2.7 extends to approximate JR solutions; the proof remains the same. We note, however, that Proposition 2.7 does not offer an efficient algorithm to find a Pareto-optimal solution that provides JR, as it is not clear how to compute Pareto improvements in polynomial time.
@@ 3 Efficiency
Our algorithm is based on a dynamic program, which iteratively considers adding new stops to the solution. Capturing our problem by a dynamic program is challenging, because each agent’s cost depends on the placement of two stops. The crucial observation that enables us to circumvent this difficulty is that, to perform cost updates in the dynamic program, it suffices to know the rightmost stop in the current solution. The relevant computation is captured in Lemma 3.1, which addresses the change in cost resulting from the addition of an extra bus stop. All missing and full proofs from this section can be found in Appendix B.
@@
Proof Sketch.
The key observation is that, after $k$ is added, the only update to their route that agent $i$ needs to consider is whether to ride the bus between $h$ and $k$ and then walk to their destination; alternatively, they can keep their original route. Their gain from riding the bus between $h$ and $k$ only depends on $h,k,\ell_{i}$, and $r_{i}$. ∎
@@
For $\alpha\in[0,1]$, we can compute a minimum-cost solution for an instance of $\alpha$-BSP in time $O(nm^{2})$.
@@
First, in the case where $b\geq 2n$, the budget is allows us to assign two dedicated bus stops to each agent. Then, for each agent we can simply establish the two best bus stops. For this, we can first find the closest bus stops to the left and to the right of their two terminals in time $O(m)$. If there is a bus stop at the agents position, we only use this bus stop. Then, we can check the at most four possible combinations of using the bus from each of the bus stops closest to the agent and compare it to the cost of walking. Performing this for each agent leads to a total running time of $O(nm)$. The union of the best bust stops yields the smallest cost in any feasible solution for each agent. Hence, it suffices to consider instances where $b<2n$, for which we will define our dynamic program.
Let $\alpha\in[0,1]$, and consider an instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of $\alpha$-BSP where $V=\{v_{1},\dots,v_{m}\}$ with $v_{1}\leq v_{2}\leq\dots\leq v_{m}$, i.e., potential stops are sorted from left to right and $v_{j}$ represents the $j$th stop. We will assume $b\leq m$, as otherwise there is an optimal-cost solution that builds a stop at each location. For a solution $S$ with $\max S=v_{h}$ and $k>h$, let $\Delta(h,k):=c(S)-c(S\cup\{v_{k}\})$ denote the total reduction in the agents’ costs from adding stop $v_{k}$ to $S$. By Lemma 3.1, we know that $\Delta(h,k)$ only depends on $v_{k}$ and $v_{h}$, and can be computed in $O(n)$ time. Hence, we can precompute $\Delta(h,k)$ for all values of $h$ and $k$ in time $O(nm^{2})$ since both $h$ and $k$ can take at most $m$ values. Let us set up a dynamic program $\mathbf{dp}[h,c]$, where
  • $h\in\{0,1,\dots,m\}$ represents the rightmost stop that has been added to the solution, where 0 means that no stop has been added yet, and
@@
That is, we consider the position of the stop $h^{\prime}$ that precedes $h$ in the solution, and evaluate the cost reduction from adding $h$ to a solution that ends with $h^{\prime}$. Clearly, the updates can be computed in time $O(m)$, using the precomputed values $\Delta(h,k)$. As we assume that $b<2n$, our table has at most $2nm$ entries, and can be filled in time $O(nm^{2})$. Hence, together with the precomputation of the values $\Delta(h,k)$ in time $O(nm^{2})$ and initializing the table in time $O(nm)$, the total running time is $O(nm^{2})$.
It is not hard to see that our dynamic program is correct; we provide a formal proof of correctness in Appendix B. Hence, the minimum cost of a feasible solution for $\mathcal{I}$ is equal to $\min_{h\in\{0,\dots,m\}}\mathbf{dp}[h,b]\text{.}$
We can efficiently extract an explicit feasible solution of minimum cost by standard techniques. ∎
@@
First, the theorem extends to more general cost functions. Consider a setting where the time to travel between two stops depends on factors other than the distance. For instance, the bus route might encompass intervals with different speed limits, or there may be hilly or curvy roads, where the bus needs to slow down. Hence, the travel costs need not be homogeneous. However, typically the travel cost of a route only depends on the costs of its segments. To capture this, we introduce the notion of separable travel costs, and formally introduce the separable-cost BSP problem, which generalizes the $\alpha$-BSP problem defined earlier in the paper.
A function $d\colon\mathbb{Q}^{2}\to\mathbb{Q}$ is called separable if for all $x<y<z$ it holds that $d(x,z)=d(x,y)+d(y,z)$. An instance $\mathcal{I}$ of separable-cost BSP is given by a tuple $\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ and separable cost functions $d^{W}:V\times V\to{\mathbb{Q}}$ and $d_{i}^{B}:V\times V\to{\mathbb{Q}}$ for all $i\in N$ (where the superscripts refer to walking and taking the bus). The agents’ costs are then defined as
@@
For separable-cost BSP, we can compute the minimum cost in time $O(nm^{2})$.
@@
For $\alpha$-BSP with existing bus stops, we can compute the minimum cost in time $O(nm^{2})$.
As a third extension, we consider the case where bus stops do not have identical costs: indeed, construction costs may vary depending on, e.g., ease of access. An instance of $\alpha$-BSP with bus stop costs consists of an instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of the base model together with a budget function $\gamma\colon V\to\mathbb{N}$. A solution $S\subseteq V$ is feasible if $\sum_{i\in S}\gamma(i)\leq b$. Clearly, we can still apply the dynamic program in Theorem 3.2. However, we can no longer assume that $b\leq m$. Hence, we obtain a running time of $O(nm^{2}+m^{3}b)$, which is only pseudo-polynomial due to the dependency on $b$. If the bus stop costs are represented in unary, the running time remains polynomial. However, for bus stop costs represented in binary, we obtain a computational hardness result, via a reduction from Knapsack.
@@
Let $\alpha\in[0,1)$. Then, the following decision problem is $\mathsf{NP}$-complete: given an instance $\mathcal{I}$ of $\alpha$-BSP with bus stop costs represented in binary and a rational number $q\in\mathbb{Q}$, decide if there exists a feasible solution $S$ with $c^{\mathcal{I}}(S)\leq q$.
We conclude this section with a structural result regarding minimum-cost solutions. Interestingly, as long as both terminals of all agents belong to the set of potential bus stops, there is a minimum-cost solution that places all stops at the agents’ terminals. We remark that this is the only place in this section where the assumption $\mathcal{A}(\mathcal{I})\subseteq V$ plays a crucial role.